OpenStudy (anonymous):

Use the quadratic formula to solve the equation. –4x2 – 3x + 2 = 0

OpenStudy (anonymous):

x= 1/8(-3- squareroot 41)

OpenStudy (anonymous):

identify your a, b, c in the quadratic equation: \(\large ax^2+bx+c=0 \) then just plug into the quadratic formula: \(\large x=\frac{-b\pm \sqrt{b^2-4ac}}{2a} \)